OpenStudy (gabylovesyou):

There are 384 nails in a 16-ounce jar. How many nails are there per ounce in the jar? 12 6,144 0.0417 24

jimthompson5910 (jim_thompson5910):

(384 nails)/(16 ounces) = (x nails)/(1 ounce) 384/16 = x/1 x = ??

OpenStudy (gabylovesyou):

why did i ask this... lol i asked the wrog question.... its 24

OpenStudy (gabylovesyou):

wrong*

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (gabylovesyou):

lol thanks

jimthompson5910 (jim_thompson5910):

np